OpenStudy (anonymous):

A survey of 1,500 high school students was given to find their favorite school subject. The pie chart below gives the results. Of the students surveyed, how many prefer Art?

OpenStudy (bahrom7893):

where's the pie chart?

OpenStudy (bahrom7893):

24.8% prefer art: 1,500 - 100% x - 24.8% x = 1,500*24.8%/100% = 15*24.8

OpenStudy (anonymous):

That's my final answer?

OpenStudy (bahrom7893):

can't u multiply? 15*24.8 = 372 <= final answer